OpenStudy (anonymous):

4+4-4*4*4+4+4-4-4+4+4/4 = ???

OpenStudy (mayankdevnani):

so confusing

OpenStudy (anonymous):

hahahaha No i don't think so !!!

OpenStudy (mayankdevnani):

-51 i got it

OpenStudy (mayankdevnani):

right @hartnn

OpenStudy (mayankdevnani):

apply the method of bodmas

OpenStudy (mayankdevnani):

soory DMAS

OpenStudy (mayankdevnani):

first divide then multiply then add and last subtract

OpenStudy (anonymous):

tell me the final answer

OpenStudy (mayankdevnani):

for more convinent, we can also written as 4+4-(4*4*4)+4+4-4-4+4+(4/4)=-51

OpenStudy (mayankdevnani):

-51 is final answer

OpenStudy (unklerhaukus):

=4+4-(4*4*4)+4+4-4-4+4+(4/4) =8-64+8-8+4+1 =8-64+5 =13-64 =-51

OpenStudy (anonymous):

Tanx guys

OpenStudy (mayankdevnani):

welcome